# Hierarchical models for small area estimation using zero-inflated forest inventory variables: comparison and implementation

### v1

@@ 2.5.3 Metrics for comparison
We use the bias metric to evaluate the bias of estimates, estimates of the RMSE, and unit-level predictions in subsequent sections. Similarly to the RMSE, the bias metric is evaluated empirically. Finally, the indicator of coverage for a given uncertainty interval is
@@ 2.6 Simulation
A simulation study was used to assess qualities of the estimators introduced in Section 2.2. Simulation studies are particularly useful in SAE research as they allow for us to assess estimators against simulated true parameter values, letting us gain accurate insight into estimator performance. In order to assess estimators in a manner that is fair and realistic, we utilize methodology introduced in White et al., 2024a for our simulation study. Briefly, the approach uses a nearest neighbors (NN) algorithm on auxiliary data weighted with bootstrap inclusion probabilities to impute forest inventory attributes at each population unit. Formally, the algorithm to generate the population for a given state is defined in White et al., 2024a Algorithm 1.
It is important to note that we generate simulated populations separately in Washington and Nevada. Further, as discussed in Section 2.1, we use tcc, elev, and ppt as auxiliary data in the NN matching in Washington, and tcc, elev, and tri in Nevada. We also use tnt as a stratification variable for generating the simulated population. These auxiliary variables are centered and scaled before the matching occurs. Without the centering and scaling step, variables of different magnitudes would hold different weight in the NN search. We implement the generation of the simulated population via the kbaabb R package (White et al., 2024b, ).
After generating the simulated population we take design-based samples from the simulated population for the purposes of estimator assessment. In order to create one sample from the simulated population, we take a simple random sample from each county of the same size as the number of FIA plots in that county. Therefore, both the county and overall sample sizes remain constant.
@@ 3 Results
We now turn to discussing the results of the simulation study and FIA data application in Sections 3.1 and 3.2, respectively. In Section 3.1, to evaluate performance of the estimators, we assess metrics introduced Section 2.5.3. The FIA data application present county-level biomass estimates and compare estimators based on cross-validation.
@@ 3.1 Simulation study
We fit and evaluated the nine estimators introduced in Section 2 across all samples taken from the simulated population generated in Section 2.6. These estimators are fit separately between the two states considered in this study, Nevada and Washington. We first turn to evaluating performance metrics in Washington, and then to Nevada. Generally, estimator performance differs between the two states, likely due to the substantial differences in the ecological landscape of the states, with Nevada’s forests primarily occupying high elevation sky islands, and Washington’s forests primarily on the west side of the Cascade mountain range.
Figure 1 displays values for the four performance metrics evaluated in Nevada, with each point on the plot represented a county-performance metric-estimator combination. Turning our attention to Figure 1a, we see that the single stage estimators exhibit the most bias. Further, we see that estimators that do not allow for county-specific variances tend to exhibit more bias than those that allow for county-specific variances. The least bias estimators are the B ZI CVI CRV and the B ZI CVI SVI CRV, both performing quite similar to each other. The CVC models, when compared to their CVI counterparts, tend to exhibit more bias after CRV and/or SVI terms are added. When turning to Figure 1b, the negative effect of the CVC effects in Nevada becomes even clearer, as we see these models exhibit much higher RMSE than those without the CVC terms. All two-stage estimators with the CVI effect perform comparably in terms of RMSE, with the others performing comparably to each other. Figure 1c displays the bias of , indicating if an estimator is over- or under-estimating its variability. The B CVI and B CVC tend to substantially under estimate their variability, likely due to the poorly specified underlying model. Somewhat surprisingly, the B ZI CVI estimator also exhibits some negative bias. The F ZI CVI does quite a good job of estimating its RMSE, but has more variability in those estimates than the B ZI CVC, B ZI CVI CRV, B ZI CVC CRV, B ZI CVI SVI CRV, and B ZI CVC SVI CRV, all of which estimate the RMSE quite well, but have a few negative outliers. Figure 1d displays the 95% uncertainty interval coverage, where we can see the B CVI and B CVC estimators performing very poorly. The F ZI CVI performs better than its Bayesian counterpart in this case, but both have some counties with very low coverage rates. The remaining estimators all perform quite well and similar to each other, with the B ZI CVI SVI CRV performing the best on median.
Figure 2 displays values for the four performance metrics evaluated in Washington, with each point on the plot represented a county-performance metric-estimator combination. Figure 2a displays the bias of each estimator, and we see in Washington that the F ZI CVI estimator exhibits the most bias, substantially more than its Bayesian counterpart. Initially this may come as a surprise due to the estimators similar structure, but one must consider the differences in the back-transformation of the response variable between these estimators. While the frequentist estimator back-transforms the predictions from each model, the Bayesian estimators back-transforms each sample from the posterior predictive distribution, potentially leading to more sensible estimates under certain conditions. The B CVI and B CVC estimators tend to exhibit greater magnitude bias than the two-stage Bayesian estimators, all of which exhibit similar amounts of bias to each other. One notable observation is the B ZI CVC CRV and B ZI CVC SVI CRV estimators exhibit a slight amount more bias than their CVI counterparts, and when we turn to Figure 2b we can see their RMSE is higher than their CVI counterparts. The B ZI CVI CRV and B ZI CVI SVI CRV estimators have the lowest RMSE, while the F ZI CVI estimator has the highest RMSE. Other two-stage Bayesian estimators exhibit very similar RMSE to each other, and the single-stage estimators exhibit a bit higher RMSE. Turning to Figure 2c, we see the bias of the RMSE estimator. Notably, the F ZI CVI and B CVI estimators exhibit the most bias here, while the other estimators exhibit very low bias. Once CRVs are accounted for, we see almost no bias. Finally, turning to Figure 2d, we examine the 95% uncertainty interval coverage rate. The B CVI and B CVC estimators have very poor coverage, while the other estimators, all of which are two-stage, have decent coverage among themselves. Notably, the estimators which account for CRVs and have CVC effects exhibit the best coverage rates.
Stepping back from the details of Figures 1 and 2, we can gain some broader insights about the estimators and the simulation study. First of all, it is striking that the estimators that include a SVI perform almost identically to their non-spatial counterparts. This is not to say that spatial effects are unuseful for estimating biomass or other forest attributes, in fact we know from other studies these effects are quite useful (Finley et al.,, 2024). Further, from the cross-validation carried out in Section 3.2 we see more substantial improvements with the spatial models. It is the case though, that the simulated population we generated does not necessarily reflect the spatial structure of the true population, and in fact upon inspection it has very little spatial structure. Incorporating some spatial smoothing into the population generation methodology would help improve the simulated population’s utility for assessing these spatial models. Another broad observation we can make from the simulation study is that the CVCs generally did not improve estimation, and in some cases they worsened estimation. We see this particularly in Nevada, where county-level biomass is more homogeneous across the state. In Washington, the estimators with CVCs did have better 95% coverage rates than their CVI counterparts, but at the cost of higher bias and RMSE. The use of CVCs may be more useful for models fit across larger spatial domains (such as the entire United States), where we would expect the effect of predictors on the response to vary substantially. States with more diverse forest types may also benefit from the use of CVCs.
@@ 3.2 FIA data application
We applied the estimators discussed to actual plot data collected by FIA in both Nevada and Washington. We first compare the estimators using 10-fold cross-validation, and then turn to discussing estimates produced by the B ZI CVI SVI CRV estimator, an estimator that showed favorability in the simulation study and in cross-validation.
We performed 10-fold cross-validation at the unit-level to help assess the estimators performance, with the unit-level predictions as a proxy for estimation of county means. Results from the cross-validation are shown in Table 3. To compute 95% uncertainty interval coverage rates (“Coverage” in Table 3), we used the quantiles of the posterior predictive distribution of for the Bayesian estimators. We omit 95% confidence interval coverage for the frequentist estimators as the task of computing closed form or bootstrap prediction intervals at the unit-level for this two-stage model is beyond the scope of this article. Regarding the results, we can see similar patterns as we saw in Figures 1 and 2, but now we can better see the performance improvement associated with the spatial models when considering the root mean square prediction error (RMSPE) and bias. We see that across states bias and coverage rates were favorable for all estimators that included a county-specific variance term, and in Nevada for all estimators that account for zero-inflation. Examining the results of this cross-validation is helpful for model evaluation and in this case provides similar insights to the simulation study, and in Section 4 we discuss the tradeoffs of the two model evaluation methods.
@@
In both Nevada and Washington, the B ZI CVI SVI CRV estimator performs quite well when considering all computed metrics in the simulation study and cross-validation, so we chose to use this estimator as an example for producing pixel-level predictions and county-level estimates in Washington and Nevada. Figure 3a, b show pixel-level estimates of biomass probability and Figure 3c, d show estimated biomass (Mg/ha) in both Washington and Nevada produced by the B ZI CVI SVI CRV estimator. While we are not particularly interested in the pixel-level predictions for the study at hand, these sorts of maps can be very useful for management at very small spatial scales, such as the stand level. In this Bayesian framework, not only do we have pixel-level predictions, but we also have pixel-level uncertainty predictions. Further, we can easily summarize these pixel level predictions to get estimates and uncertainty estimates for any area that may be of interest. For the study at hand, we only produce estimates for counties, but we could have just as easily produced estimates for any custom small area of interest (e.g., ecoregions, watersheds, fires, stands, etc.) with any of the Bayesian estimators.
Figure 4 shows the estimated biomass (Mg/ha) in both Nevada and Washington produced by the B ZI CVI SVI CRV estimator. These sorts of estimates are what we are most interested for SAE, and for demonstration we only produce estimates at the county-level. County-level estimates of biomass show the stark change in biomass across the cascades in Washington state, and the relatively constant county-level biomass across the state of Nevada.
4 Discussion
@@
Our work implemented and compared nine model-based approaches from both the frequentist and Bayesian statistical paradigms to estimating biomass in counties in the states of Nevada and Washington. The Bayesian estimators used allow for much flexibility in model specification, allowing for CVIs, CVCs, CRVs, and SVIs. We assessed the proposed estimators through a simulation study using design-based samples from a simulated population generated by a bootstrap-weighted NN technique, and through cross-validation at the unit-level on FIA data. Results suggest that accounting for zero inflation through a two-stage approach, the inclusion of CRVs, and the inclusion SVIs yield the best performing estimators.
It is an inherently difficult task to compare estimators for the purposes of SAE given to the small sample size in areas of interest and lack of true parameter values at the granularity of the area of interest. We implemented two methods for evaluating the proposed estimators, each of which have tradeoffs. First, we generated a simulated population for estimation evaluation. This approach allows for us to assess estimators empirically by sampling repeatedly from the simulated population and comparing estimates to the true value for the response in each county. These assessments are only insightful in the case that (1) the simulated population does not unfairly portray the estimators by being generated from some method too similar to the proposed estimators; and (2) the simulated population is similar enough to the true population that fitting these estimators to samples from the simulated population is representative of fitting the estimators to a sample of the true population. We achieve (1) here by using the bootstrap-weighted NN simulated population generation technique proposed by White et al., 2024a , but partially miss the mark regarding (2) as the simulated population does not have similar spatial structure to the true population, impeding our ability to accurately assess the models with a space varying intercept. Second, we used cross-validation to assess prediction made onto unit-level observations. This approach does not require assumptions about how the population is generated or otherwise, but only allows for assessment of unit-level predictions. In our case, the cross-validation analysis yielded very similar results to the simulation study, and we were able to more clearly see the benefit of the spatial models. However, cross-validation is only possible for unit-level small area estimators and we must tolerate unit-level predictions as a proxy for small area estimates.
Our analyses provide insight as to how we may want to go about estimating biomass, or other continuous zero-inflated forest attributes, in areas of mixed landscape types and how a practitioner may want to assess their modeling choices. Given our changing climate and increasingly prolific disturbances, such as fires, producing accurate estimates of forest attributes in small areas that are zero-inflated is of the upmost importance understanding our ever-changing landscape. In future work, we hope to test the proposed estimators across a variety of regions such as recently burned areas, forest stands, and other areas with great ecological importance. Further, we hope to improve our methodology for creating simulated populations in order to better assess the quality of spatial area- and unit-level small area estimators. We also hope to consider temporal effects in these zero-inflated estimators, to be able to estimate change in biomass across time scales.
@@
Acknowledgments
The authors thank the Forest Inventory & Analysis Program for the data.
@@ Acknowledgments
This work was supported by the USDA Forest Service Forest Inventory and Analysis Program, Region 9, Forest Health Protection, Michigan State University AgBioResearch, National Science Foundation DEB-2213565 and DEB-1946007, and National Aeronautics and Space Administration CMS Hayes-2023. The findings and conclusions in this publication are those of the authors and should not be construed to represent any official US Department of Agriculture or US Government determination or policy.

### v2

@@ 2.6.3 Metrics for comparison
where is computed empirically across the design-based samples. We use the bias metric to evaluate the bias of estimates, estimates of the RMSE, and unit-level predictions in subsequent sections. Similar to the RMSE, the bias metric is evaluated empirically. Finally, the indicator of coverage for a given uncertainty interval is
@@ 2.7 Simulation
We conducted a simulation study to evaluate the performance of the estimators introduced in Section 2.3. Simulation studies are especially valuable in SAE assessment, as they allow for comparisons against known true parameter values, thereby providing direct insight into estimator performance. To ensure a fair and realistic evaluation, we followed the methodology proposed by White et al., 2024a .
Briefly, this approach uses a -nearest neighbors (NN) algorithm applied to auxiliary data, with neighbors weighted by bootstrap inclusion probabilities, to impute forest inventory attributes at each population unit. The complete algorithm for generating a population for a given state is presented in Algorithm 1 of White et al., 2024a .
Simulated populations were generated separately for Washington and Nevada. As described in Section 2.1, auxiliary variables used in the NN matching included tcc, elev, and ppt in Washington, and tcc, elev, and tri in Nevada. In both states, tnt was used as a stratification variable to distinguish between treed and non-treed areas. Prior to matching, all auxiliary variables were centered (mean zero) and scaled (unit variance) to prevent differences in variable magnitude from disproportionately influencing the NN search. The population generation process was implemented using the kbaabb R package (White et al., 2024b, ).
Following population generation, we drew design-based samples to assess estimator performance. For each simulation replicate, we selected a simple random sample within each county, matching the number of FIA plots observed in that county. This ensured that both county-level and overall sample sizes remained consistent with the original FIA data.
@@ 3 Results
This section presents the results of the simulation study and the FIA data application, discussed in Sections 3.1 and 3.2, respectively. In Section 3.1, we assess estimator performance using the metrics defined in Section 2.6.3. Section 3.2 provides county-level biomass estimates and evaluates estimator performance based on cross-validation.
@@ 3.1 Simulation Study
We evaluated the nine estimators introduced in Section 2 using repeated samples drawn from the simulated populations described in Section 2.7. Model fitting was performed separately for Nevada and Washington to reflect state-specific differences in data-generating processes. Estimator performance was assessed using the metrics defined in Section 2.6.3. Notably, performance varied between states, likely due to substantial ecological differences: Nevada’s forests are largely confined to high-elevation sky islands, whereas Washington’s forests are primarily located west of the Cascade Range, where precipitation and productivity are substantially higher.
Figure 1 presents the estimator performance metrics for Nevada. Each point in the figure represents a county–metric–estimator combination. In Figure 1a, we observe that single-stage estimators tend to exhibit the greatest empirical bias. Additionally, estimators that do not account for county-specific variances generally perform worse in terms of bias than those that do. The least biased estimators are the B ZI CVI CRV and B ZI CVI SVI CRV, which yield nearly identical results. Comparisons between the CVC and CVI model structures reveal that, in Nevada, the inclusion of CVC terms typically increases empirical bias once CRV and/or SVI terms are added.
These trends are even more pronounced in Figure 1b, which displays empirical root mean squared error (RMSE). Models with CVC terms consistently show higher RMSE values than their CVI-based counterparts. Among the CVI-based two-stage estimators, RMSE performance is comparable across models. Conversely, models without CVI terms also perform similarly to one another but lag behind the CVI-based estimators.
Figure 1c shows the empirical bias in RMSE estimation, which reflects whether an estimator under- or over-estimates its own uncertainty. The B CVI and B CVC models significantly underestimate their RMSE, likely due to poor model specification. Somewhat unexpectedly, the B ZI CVI estimator also shows RMSE underestimation, despite accounting for zero-inflation. The F ZI CVI model estimates its RMSE reasonably well but displays greater variability compared to the B ZI models that include CRV and/or SVI effects, which generally yield accurate RMSE estimates but occasionally produce negative outliers.
Lastly, Figure 1d illustrates empirical 95% uncertainty interval coverage. The B CVI and B CVC estimators again perform poorly, exhibiting undercoverage across many counties. While the F ZI CVI estimator performs better than its Bayesian counterpart in this regard, both show some counties with notably low coverage. The remaining estimators—including those incorporating zero-inflation, CRV, and SVI components—demonstrate strong and consistent coverage performance, with the B ZI CVI SVI CRV estimator yielding median coverage closest to the desired value, 0.95.
Figure 2 presents the estimator performance metrics for Washington, where each point represents a county–metric–estimator combination. Figure 2a displays empirical bias. The F ZI CVI estimator exhibits the highest bias among all estimators—substantially more than its Bayesian counterpart, B ZI CVI. Although the two estimators are structurally similar, this discrepancy is likely attributable to differences in how each handles the back-transformation of the response variable. The frequentist estimator applies the back-transformation directly to the predicted mean, whereas Bayesian estimators apply it to each sample from the posterior predictive distribution, which can yield more accurate and robust estimates, especially in non-linear settings.
The B CVI and B CVC estimators also exhibit relatively high bias compared to the two-stage Bayesian models, all of which show comparable and generally lower bias. Interestingly, the B ZI CVC CRV and B ZI CVC SVI CRV estimators display slightly higher bias than their CVI-based counterparts. As shown in Figure 2b, these same estimators also have higher empirical RMSE. The B ZI CVI CRV and B ZI CVI SVI CRV estimators yield the lowest RMSE values overall, while the F ZI CVI estimator exhibits the highest RMSE. Other two-stage Bayesian estimators have similar RMSE values, with single-stage estimators performing slightly worse.
Figure 2c shows the empirical bias in RMSE estimation. The F ZI CVI and B CVI estimators perform poorly when estimating their own variability, whereas most other estimators—particularly those incorporating CRV terms—demonstrate minimal bias and precise estimation of their own variability. This suggests that accounting for county random variation improves the accuracy of uncertainty estimation.
Finally, Figure 2d presents empirical 95% uncertainty interval coverage. The B CVI and B CVC estimators perform poorly in this regard, with consistently low coverage rates across counties. In contrast, the remaining estimators—all two-stage models—achieve more reliable coverage. Notably, those incorporating both CVC effects and CRV terms, such as B ZI CVC SVI CRV, exhibit the best coverage performance.
Stepping back from the specific results presented in Figures 1 and 2, several broader insights emerge regarding the estimators and the simulation study design. One notable finding is the minimal performance difference between estimators that include a spatially varying intercept (SVI) and their non-spatial counterparts. This observation should not be interpreted as evidence against the utility of spatial effects in modeling forest attributes. On the contrary, previous studies have demonstrated the substantial benefits of incorporating spatial structure in such settings (Finley et al.,, 2024). Moreover, the cross-validation results presented in Section 3.2 show that spatial models often yield more accurate predictions when applied to real FIA data.
The lack of benefit from SVI models in the simulation study is likely attributable to the limited spatial structure in the simulated populations. Upon inspection, the populations generated via the NN-based approach exhibit weak spatial autocorrelation. This suggests that the simulation framework, while useful for certain evaluations, may not be well suited for assessing the potential advantages of spatial models. Future simulation designs could benefit from incorporating spatial smoothing or explicitly modeling spatial dependence during population generation to more accurately reflect the structure observed in real-world forest systems.
Another general finding is the limited utility of models incorporating CVC effects. In many cases—particularly in Nevada—CVC models not only failed to improve estimator performance but also introduced greater bias and RMSE. This may be due to the relative homogeneity of county-level biomass in Nevada, which provides limited justification for allowing predictor effects to vary by county. In contrast, in Washington—where forest types and biomass distributions are more heterogeneous—CVC-based estimators exhibited better 95% uncertainty interval coverage than their constant-coefficient counterparts. However, these gains came at the cost of increased bias and RMSE. These results suggest that CVC effects may be more beneficial in broader applications involving larger spatial domains—such as national-level analyses—or in regions with more diverse ecological conditions, where predictor–response relationships are expected to vary more substantially across space.
@@ 3.2 FIA data application
We applied the candidate estimators to FIA plot data collected in Nevada and Washington. We first evaluated performance using 10-fold cross-validation, followed by an examination of county-level biomass estimates produced by the B ZI CVI SVI CRV estimator, which performed favorably in both the simulation study and cross-validation.
Cross-validation was conducted at the unit level, with unit-level predictions used as a proxy for estimating county-level means. Results are presented in Table 3. For Bayesian estimators, 95% uncertainty interval coverage rates (“Coverage” in Table 3) were computed using quantiles of the posterior predictive distribution of . Coverage is not reported for the frequentist estimator, as deriving either closed-form or bootstrap-based prediction intervals at the unit level for the two-stage model is beyond the scope of this study.
The results exhibit patterns similar to those observed in the simulation study (Figures 1 and 2), but with clearer improvements associated with spatial models—particularly in terms of root mean square prediction error (RMSPE) and empirical bias. Across both states, estimators incorporating a county-specific variance term consistently showed favorable bias and coverage rates. In Nevada, estimators accounting for zero-inflation also performed notably well.
These cross-validation results reinforce findings from the simulation study and provide complementary insights into estimator performance under realistic sampling conditions. We further discuss the relative merits of simulation-based and cross-validation-based evaluation strategies in Section 4.
@@
In both Nevada and Washington, the B ZI CVI SVI CRV estimator performed consistently well across all metrics evaluated in the simulation study and cross-validation. As a result, we selected this estimator to produce pixel-level predictions and county-level biomass estimates in each state.
Figure 3a and 3b display pixel-level estimates of biomass probability, while Figure 3c and 3d show pixel-level estimates of biomass (Mg/ha) in Washington and Nevada, respectively. Although pixel-level predictions are not the primary focus of this study, such maps can be valuable for forest management at finer spatial scales, such as the stand level. Within the Bayesian framework, pixel-level predictions are accompanied by corresponding uncertainty estimates, enabling flexible spatial summarization. While we report aggregated estimates at the county level for this analysis, the same framework allows for the generation of estimates and associated uncertainties for any user-defined spatial unit (e.g., ecoregions, watersheds, fire perimeters, or management units) using any of the Bayesian estimators considered.
Figure 4 presents the resulting county-level biomass estimates for both states, generated using the B ZI CVI SVI CRV estimator. These county-level estimates represent the primary inferential target for SAE in this study. As expected, the results for Washington show a pronounced gradient in biomass across the Cascade Range, while biomass estimates in Nevada appear relatively homogeneous across counties.
4 Discussion and Conclusion
@@
We implemented and compared nine model-based estimators—spanning both frequentist and Bayesian frameworks—for estimating biomass at the county level in Nevada and Washington. The Bayesian estimators offered flexibility in model specification, allowing the inclusion of CVIs, CVCs, CRVs, and SVIs. To evaluate these estimators, we used two complementary strategies: a simulation study based on design-based sampling from a bootstrap-weighted NN-generated population, and unit-level cross-validation using FIA data. Overall, the results indicate that estimators incorporating two-stage zero-inflation modeling, CRVs, and SVIs performed best.
One notable finding emerged from a comparison between the F ZI CVI and B ZI CVI estimators in Washington. While these estimators yielded similar results in Nevada, they diverged considerably in Washington. Given their comparable model forms, such divergence is initially surprising. However, the discrepancy stems from differences in how the response variable is back-transformed in the two paradigms. In the frequentist setting, predictions are back-transformed using the maximum likelihood point estimate , whereas the Bayesian approach applies the back-transformation to each draw from the posterior predictive distribution at location . In Washington—where the biomass distribution is heavily skewed—the frequentist approach fails to capture the non-symmetric distributional shape, resulting in biased predictions. The Bayesian approach, by contrast, preserves the skewness and produces more realistic estimates.
Evaluating estimators for SAE is inherently challenging due to small sample sizes within target areas and the absence of true parameter values at those granular scales. To address this, we adopted two evaluation frameworks, each with tradeoffs. The first—simulation-based evaluation—enables empirical comparison against known true values through repeated sampling from a synthetic population. This method is valid provided two conditions are met: (1) the simulated population is not overly similar to the estimation models, and (2) it sufficiently resembles the real-world population. We addressed (1) by adopting the bootstrap-weighted NN generation method proposed by White et al., 2024a . However, condition (2) was only partially met. Specifically, the generated population exhibited limited spatial autocorrelation—unlike what is typically observed in real forest systems. As a result, estimators incorporating SVIs did not demonstrate their full potential in the simulation setting.
The cross-validation analysis offered a complementary perspective, free from assumptions about the population generation process. While it does not directly assess small area means, it allows for meaningful evaluation of unit-level prediction accuracy. In our case, the cross-validation results aligned well with those from the simulation study but also highlighted the added value of spatial models—particularly those including SVIs and CRVs—under realistic conditions. A limitation of this approach is that it is only applicable to unit-level models and necessitates accepting unit-level predictions as a proxy for small area estimates.
Our findings offer practical guidance for estimating biomass—or other continuous, zero-inflated forest attributes—across heterogeneous landscapes. In the context of a changing climate and increasing disturbance regimes such as wildfires, the need for accurate, small area estimates of forest structure is more critical than ever. One limitation of this study is the lack of spatial structure in the simulated populations, which limited the utility of simulations for evaluating spatial estimators. To address this, future work will incorporate spatial dependence into population generation using non-parametric spatial smoothing techniques.
Beyond methodological refinement, future research will focus on extending these estimators to additional spatial units of ecological interest—such as recently burned areas, forest stands, or watersheds—and to temporal applications for monitoring biomass change over time. These directions will enhance the utility of model-based SAE methods for forest inventory and broader environmental monitoring.
@@
Competing interests statement
JKY is employed by RedCastle Resources, Inc. GWW is part-time employed by RedCastle Resources, Inc. The remaining authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
@@
This work was supported by the: US Forest Service (USFS) Forest Inventory and Analysis Program; USFS Region 9, Forest Health Protection; Michigan State University AgBioResearch; National Science Foundation DEB-2213565 and DEB-1946007; National Aeronautics and Space Administration CMS Hayes-2023. The findings and conclusions in this publication are those of the authors and should not be construed to represent any official US Department of Agriculture or US Government determination or policy.